OpenStudy (anonymous):

Esmeralda is graphing a polynomial function as a parabola. Before she begins graphing it, explain how to find the vertex. Make sure you include how to determine if it will be a maximum or minimum point. Use an example quadratic function to help you explain and provide its graph.

OpenStudy (anonymous):

@chrissyC.

OpenStudy (anonymous):

If the equation is y = 3(x + 4)2 - 6, the value of h is -4, and k is -6. To convert a quadratic from y = ax2 + bx + c form to vertex form, y = a(x - h)2+ k, you use the process of completing the square. Let's see an example. Convert y = 2x2 - 4x + 5 into vertex form, and state the vertex.

OpenStudy (anonymous):

lol ok thx peanut

OpenStudy (anonymous):

anytime Cx i doubt it helped anyways

OpenStudy (anonymous):

lol it did your so helpful

OpenStudy (anonymous):

Haha yay xD.

OpenStudy (anonymous):

ok i got one more peanut you up for it

OpenStudy (anonymous):

uh okai ;-;

OpenStudy (anonymous):

Heinz boasts that he can predict the degree of a polynomial function just by looking at the end behavior. Can Heinz do this? Explain.

OpenStudy (anonymous):

No, you can't predict the degree - only whether the highest power of x is odd or even.

OpenStudy (anonymous):

ok tytytytyty sooooooo much peanut

OpenStudy (anonymous):

lml yw jelly <3

OpenStudy (anonymous):

<3

OpenStudy (anonymous):

=^.^=